OpenStudy (anonymous):

(x-6) divided by 5 plus 2 divided by 8 = (2x-7) divided by 4

OpenStudy (studygurl14):

\(\Large\frac{x-6}{5}+\frac{2}{8}=\frac{2x-7}{4}\)

OpenStudy (studygurl14):

First step, simplify the \(\large\frac{2}{8}\)

OpenStudy (kendricklamar2014):

@SilverPersia